A convertible preferred shares is priced at R100, with a conversion ratio of five. If the common shares move to R25 per share, what would the preferred shareholder’s return be if they converted now?
Answer #1

Answer:

Convertible Preferred Share : R 100

Common Share : R 25

Conversion Ratio is 5, that means 5 Common Shares for 1 Preferred Share

Value of Common Shares after Conversion : R25*5= R125

Return of Preferred Shareholder if converted : R125-R100= R 25

% Return : 25*100/100= 25%